OpenStudy (anonymous):

The set of real x for which f(g(x)) = g(f(x)), where f(x) = x + 2 and g(x) = 3/x is ?

OpenStudy (amistre64):

you want inverses eh....

OpenStudy (amistre64):

or does this look more like a "system" of equations ??

OpenStudy (amistre64):

f(g(x)) = (3/x) +2 g(f(x)) = 3/(x+2)

OpenStudy (amistre64):

(3+2x)/x = 3/(x+2) (3+2x)(x+2) = 3x 3x +6 +2x^2 +4 = 3x 2x^2 +10 = 0 2(x^2 +5) = 0 x^2 + 5 = 0 x^2 = -5 aint possible no real solutions....

OpenStudy (amistre64):

should been: 3x +6 +2x^2 +4x = 3x forgot the "x" on the 4 2x^2 +4x +6 = 0 x^2 +2x +3 = 0 still aint got no "real" values....

OpenStudy (anonymous):

thank you so much!